OpenStudy (anonymous):

what times what gives u -60 but adds up to -11

OpenStudy (johnweldon1993):

\[\large x \times y = -60\] \[\large x + y = -11\] Substitution method can be used here...

OpenStudy (anonymous):

\[a \times b = -60\] \[a + b = -11\]

OpenStudy (anonymous):

-15 and 4